OpenStudy (imtiaz7):

If the frequency of a pendulum is four times greater on an unknown planet then it is on earth, then the gravitaional constant on that planet is: a) 16 times greater b) 4 times greater c) 4 times lower d) 16 times lower e) 24 times lower Give the correct answer and explain it. @sarah786 @2Angel4 @Muskan @jacalneaila

OpenStudy (kittiwitti1):

@AllTehMaffs

OpenStudy (anonymous):

The restoring force of a pendulum \[-mg \sin \theta = ma\] For small angles, \[\sin \theta = \theta\] so \[a=\frac{\mathrm{d}^2v}{\mathrm{d}t^2}=\ell \frac{\mathrm{d}^2\theta}{\mathrm{d}t^2}=-g\theta\] giving the simple diff eq \[\frac{\mathrm{d}^2\theta}{\mathrm{d}t^2}+\frac{g}{\ell}\theta=0\] with the solution \[\theta=\theta_{max} \sin(\omega t)\] \[ \text{angular frequency} \ = \omega = \sqrt{\frac{g}{\ell}}\] Earth \[ \omega_{E}= \sqrt{\frac{g_{E}}{\ell}}\] If \[ \omega_{p} = 4\omega_{E}\] then \[ 4\omega_{E}= \sqrt{\frac{g_{p}}{\ell}}\] \[4\sqrt{\frac{g_{E}}{\ell}}=\sqrt{\frac{g_{p}}{\ell}}\] bring the 4 into the radical to solve

OpenStudy (anonymous):

\[\sqrt{\frac{16g_{E}}{\ell}}=\sqrt{\frac{g_{p}}{\ell}}\] \[16g_{E}=g_p\]

OpenStudy (anonymous):

Great Master Maff @AllTehMaffs :) ^^